OpenStudy (starriebaby):

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles? Enter your answer in the box. inches Part B How many miles are represented by 1 4 inch on the map? Enter your answer in the box. miles Save

OpenStudy (jonathan34):

i hate u

OpenStudy (jonathan34):

sorry wrong person kind off same pic lol

Mehek (mehek14):

For Part A, lets set up proportions \(\dfrac{2}{25}=\dfrac{x}{60}\) Can you cross multiply that for me?

OpenStudy (starriebaby):

ok i will

OpenStudy (starriebaby):

x over 120 or the other way around

Mehek (mehek14):

You wold multiply 2 * 60 and 25 * x

OpenStudy (starriebaby):

yes is my answer at the top right

Mehek (mehek14):

2 * 60 = 120 25 * x = 25x Divide 120/25 = x

OpenStudy (starriebaby):

4.8

Mehek (mehek14):

So that's the answer for Part A

OpenStudy (starriebaby):

ok thanks now can you help me with part b

OpenStudy (starriebaby):

or can you not see it properly i will put it back up

Mehek (mehek14):

First simplify 25/2

OpenStudy (starriebaby):

12.5?

Mehek (mehek14):

Ok so for one inch, it's 12.5 we need \(\dfrac{1}{4}\) So divide that by 4 12.5/4 =

OpenStudy (starriebaby):

ok so it is 3.125

Mehek (mehek14):

Yes

OpenStudy (starriebaby):

is that the answer and i have another question do you think you can help me with that one

Mehek (mehek14):

Yes it is.

OpenStudy (starriebaby):

could u help me

Mehek (mehek14):

Yeah

OpenStudy (starriebaby):

k i am going to take this down and post another one

OpenStudy (starriebaby):

Question 2: A circular picture is 8 inches in diameter. Part A What is the area of the picture in square inches? A. 4π square inches B. 8π square inches C. 16π square inches D. 32π square inches Part B A frame that is 2 inches wide surrounds the picture. What is the total area of the picture and the frame in square inches? A. 4π square inches B. 12π square inches C. 36π square inches D. 40π square inches Save